Please add the process!

5. 500 in/h m/s 10. 50 m/s2km/h2
Answer #1

5.)

Since 1 inch = 0.0254 m

and 1 hour = 60*60 sec = 3600 sec.

So, 500 in/h = (500 inch/h)(0.0254 m/1 inch)(1 hr/3600 sec.)

500 in/h = 0.00353 m/sec.

500 in/h = 3.53*10^3 m/sec.

9.)

Since  1 \mum = 10^-4 cm

So, 35 \mum^3 = 35*(10^-4 cm)^3

35 \mum^3 = 35*10^-12 cm^3

10.)

Since 1 km = 1000 m

and 1 hr = 3600 s

So, 50 m/s^2 = (50 m/s^2)(1 km/1000 m)(3600 s/hr)^2

50 m/s^2 = 648000 km/h^2

50 m/s^2 = 6.48*10^5 km/h

Please upvote.

Comment below if you have any query.